OpenStudy (anonymous):

ln(x+2)-ln(5x)=(ln(x+2)/(ln(5x)))is the statement true or false?

OpenStudy (tkhunny):

Very bad. \(\ln(x+2) - ln(5x) = \ln\left(\dfrac{x+2}{5x}\right)\) \(\dfrac{\ln(x+2)}{\ln(5x)} = \ln_{5x}(x+2)\) Not very related.

OpenStudy (solomonzelman):

if you remove the second ln from the right side it will be true.

OpenStudy (solomonzelman):

you are still waiting here:?\ (At least based on the posts in this thread) is your statement false or true?